the personnel department of a company has records which show the following analysis of its 200 engineers.



































age bachelor degree only master degree total
under 30 90 10 100
30 to 40 20 30

50


over 40 40 10 50
total 150 50 200

if one engineer is selected at random from  the company,find;


a)the probability that he has a bachelor degree.


b)the probabilty that he has a master degree,given that he is over 40.


c)the probabilty that he is under 30.given that he has only a bachelor degree.

Grade:

3 Answers

Yash Melkeri
12 years ago

1)150c1/200c1=150/200=3/4

2)10c1/200c1=10/200=1/20

3)90c1/200c1=90/200=9/20

adisu yalemshet
2 years ago
Suppose that a couple will have three children. Letting B denote a boy and G denote a girl. a) Draw a tree diagram depicting the sample space outcome for this experiment. b) List the sample space outcomes that correspond to each of the following events. I. All three children will have the same gender II. Exactly two of the three children will be girls. III. Exactly one of the three children will be a girl. IV. None of the tree children will be a girl. 
tekle ayteneffesu
2 years ago
the personnel departement of a company has records which show the following analyse of its 200 accountant